A bullet of mass 10 g traveling horizontally with a velocity of 300 ms- strikes a block of wood of mass 290 g which rests on a rough horizontal floor. After impact, the block and bullet move together and come to rest when the block has traveled a distance of 15 m. Calculate the coefficient of sliding friction between the block and the floor.
1
Expert's answer
2020-11-30T14:59:28-0500

Initial speed of bullet before strike =300 m/sec and Block is at rest.Conservation of momentum before and after collision:

"(10\times300 + 290\times 0)\times 10^{-3}=(10+290)\times 10^{-3}\times U"

Where U is the block-bullet system just after the strike.

"U=10 m\/esc"

Now, equation of motion after the strike.

"F_r=M a\\
-\mu R=Ma\\
-\mu (M\times g)=Ma\\
-\mu g=a......Eq[1]"

Finally system get stopped so final speed is zero and before it travelled S=15 meter.

"V^2=U^2+2as\\
0^2=10^2-2\mu g\times15\\
\mu=\cfrac{10}{3g}\\"

Using "g=10 m\/sec^2" ,

"\mu=\cfrac{1}{3}"

Hence Coefficient of friction is "\cfrac{1}{3}" .